OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. (Do not type the variable in the blank.) 4x2 + ___x + 49

OpenStudy (triciaal):

(a + b )^2 = a^2 + 2ab + b^2 equate this with 4x2 + ___x + 49 a^2 = 4x^2 and b^2 = 49 solve for a and b now write 2ab in the blank

OpenStudy (anonymous):

2ab